The difference between two numbers is ‒2. The sum of the two numbers is 18. Find the numbers.
I set up the equation as:
x-y=-2
x+y=18
I then isolated x by canceling out y:
x=-2
x=18
x=16
then I tried to plug in 16 to the x variable to get y:
x-y=-2 x+y=18
16-y=-2 16+y=18
-16 -16 -16 -16
y=-18 y=2
What am i doing wrong?

Answer by Alan3354(69443)

The difference between two numbers is ‒2. The sum of the two numbers is 18. Find the numbers.
I set up the equation as:
x-y=-2
x+y=18
I then isolated x by canceling out y:
x=-2
x=18
x=16
then I tried to plug in 16 to the x variable to get y:
x-y=-2 x+y=18
16-y=-2 16+y=18
-16 -16 -16 -16
y=-18 y=2
What am i doing wrong?
=========================
You can't just "cancel out the y." Where did you get that idea?
Then, you get x = -2 (which is wrong)
Then, x = 18 ?????
Then, x = 16 ?
================
x-y=-2
x+y=18
------------------ Add them
2x = 16
x = 8
y = 10